OpenStudy (anonymous):

write an equation in point slope form of the line that passes through (6,4) (2,1)

OpenStudy (anonymous):

slope of line passing through (a,b) and (c,d) is (d-b)/(c-a) Point slope form is y-b= m (x-a)

OpenStudy (anonymous):

that should be y-4 = 3/4 (x-6) 4y-16=3x-12 3x-4y+4=0